OpenStudy (anonymous):

which choice converts this standard parabolic equation to its simplified form? y=2(x-(-4))^2+(-21)

OpenStudy (jdoe0001):

expand the expression to find out :) keep in mind \(\bf y=2(x-(-4))^2+(-21)\implies y=2(x+4)^2+(-21) \\ \quad \\ \implies y=2(x+4)(x+4)+(-21)\)